OpenStudy (anonymous):

Evaluate: 3^2 + (5 - 2) · 4 - 6 over 3

OpenStudy (eujc21):

\[\[(3^2 + (5-2) * (4-6))/3 = (9 +( 3*-2))/3 = (9-6)/3 = 3/3 =1\]\]

OpenStudy (eujc21):

Is that the correct format?